Beam Test Meeting Mon. Mar. 19, 2012

From GlueXWiki
Jump to: navigation, search


- Hall B will take 4 pass photon beam today through Wednesday Morning

- On Wednesday, they will go to 5-pass photon beam

- On Monday March 26th, they will go to electron running for a few days

- They will begin the transition out of electron running and begin installing their new target on April 2

- The DAQ has been idle waiting for the Hall B energy calibration running, it should happen in the next few days

- The BCAL is finishing the electronics installation, it should be complete in the next few days

- The planned installation is for the electron running / target swap scheduled beginning on Monday April 2

- The temp DAQ should be working sometime this week in 126

- Online GUIs for HV, LV, and the scalars are ready